OpenStudy (anonymous):

When Sean left for college, his parents decided to move from their two thousand four hundred square-foot home into a smaller house. If the size of their house decreased by thirty percent, how many square feet is their new house

OpenStudy (anonymous):

2400 - 30(2400)/100 = 2400 - 30(24) = 2400 - 720 = 1680 Double-c hecking the answer, 10% is easily 240 70% is their new house and 7 * 240 = 1680